Vicki is 1/3 as old as her cousin, chris, plus 5 years. Vicki is 14 years old. how old is chris?
Answer by cathieb1225(41)

Chris's age is x; Vicki's age is 1/3x+5
The equation would be 1/3x+5=14
1/3x+5=14
Subtract 5 from both sides of the equation.
1/3x+5-5=14-5
1/3x=9
Divide both sides by 1/3.
1/3x/1/3=9/1/3
Since division by a fraction is the same thing as multiplying by the reciprocal:
(1/3x)(3)=(9)(3)
x=27
Chris is 27.
To check your answer, just substitute 27 for x.
1/3(27)+5=14
27/3+5=14
9+5=14
14=14